# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ III.5 Dependencies of breakdown rate
The breakdown rate of a high-field system exhibits a very strong dependence on applied field level. In both pulsed RF and DC systems this dependence is typically . This has been reported for both [58] for radio frequency and [157] pulsed dc systems. Example of breakdown rate vs. gradient measurements are shown in Figs. 21 and 22.
@@
The breakdown rate also displays a dependence on pulse length. In the traveling wave RF accelerating structures tested in the linear collider programs, measured breakdown rates typically have shown a dependency of where is the pulse length, as for example reported in [58]. An example where this dependency can be observed is by comparing the shapes of the conditioning curves shown in Figs. 17 and 18. In both cases conditioning has been carried out at successively longer pulse lengths. The change of pulse length can be clearly seen in Fig. 17. In 18 the electric field is normalized by , the conditioning curve becomes smooth showing that the pulse length scaling is correct. It is important to note that these pulse lengths for the accelerating structure tests ranged from about 50 to 300 ns, with the minimum time give by the filling time of the structures (tests were rarely carried out below the filling time) and the maximum time is given by the power limit of the klystron-based power sources driving the structures. Breakdown rate as a function of pulse length measurements have also been carried out using the pulsed DC system, but here pulse lengths were in the range of around 500 ns up to 1 ms. In this system and pulse length range pulse the length dependence was very weak or indiscernible 23. From these two sets of measurements, it appears that the pulse length dependence saturates at around 500 ns. Pulse length dependence is discussed further in IV.5.
@@
An important insight into the character of the breakdown mechanism has been found when combining the observation that the breakdown timing distribution is flat, as discussed in Subsec.III.3 and shown in 16 even when a strong pulse length dependence is observed in RF tests. Since the breakdown rate dependence is not produced by a progressively greater density of breakdowns at the end of the pulse, the conclusion must be that the system displays a memory of the pulse length of previous pulses. This is analogous to a fatigue process where a material will tend to fail after a certain number of stress and de-stress cycles, but fails at a random moment within a cycle. This memory effect is an important motivation for the dislocation-based breakdown initiation model described in Sec. IV.
@@
Another observed dependence of breakdown rate is on temperature. For this dependence the breakdown rate, for a given field level, goes down significantly when a high-field system is cooled to cryogenic temperatures. Alternatively one can say that system supports a higher field for a given breakdown rate when cooled. This type of measurement has required specialized versions of the high-field setups described in Sec. II. An early high-field measurement of the effect of cooling on field was carried out in the CTF2 facility using a beam-driven cavity cooled to liquid nitrogen temperature [19]. This experiment did not show any temperature dependence, however only the maximum field holding level was measured, that is the field level where the breakdown rate approaches 1. We will return to this point below.
@@
A subsequent RF experiment was carried using a cryogenically cooled cavity was carried out as described in [21], in which the effect of temperature on field and breakdown rate was measure at low, range, breakdown rates. The experiment used an RF test stand similar to the type described in Sec. II but where the cavity was operated at cyrogenic temperatures using a cryocooler. When cooled to 45∘K, the cavity showed a very clear, approximately 20, increase of field level for a breakdown rate of approximately . A plot of the effect of temperature on gradient is shown in Fig. 24. The publication also reports that field emission at the increased field levels enabled by the cooling was strong enough to result in a measurable dynamic decrease of cavity . We will return to field emission at cryogenic temperatures in SubSec. III.7.
@@
A similar increase of field for a given breakdown rate has also been observed in pulsed DC experiments [61]. In this case, the experimental setup consisted of a pulsed DC system similar to that described in II but where the electrodes were cooled to cryogenic temperatures. Here also the field increased by approximately 20 for a normalized breakdown rate when cooled to a temperature of around 60 ∘K and conditioned, with further improvement as the system was cooled to around 30 ∘K. Data showing the increase of field for a normalized breakdown rate in the pulsed DC system is shown in Fig. 25. These improvements are not retained upon final warm-up back to room temperature, where the behavior is roughly identical as to before cool-down. The dc systems also showed enhanced field emission at the elevated field levels, and we will return to this in Subsec. III.7.
Both the vacancy [127] and dislocation [43] models of the breakdown rate dependence on field predict very accurately the observed dependence in the radio frequency and pulsed dc measurements described above. This can be seen by the accuracy of the fit lines in Fig. 25. The accuracy of the dislocation model is shown in Fig. 39. That the models predict both field and temperature dependence is very compelling.
@@
Another dependence of breakdown rate that has been observed is on repetition rate and more generally, the time interval between pulses. The dependence on breakdown rate has been observed in measurements made in the pulsed DC system described in Sec. II.3. In an experiment described in [150], the breakdown rate was measured as the repetition rate was varied from 10 Hz to 6 kHz and is shown in Fig. 27. The breakdown rate clearly goes down as the repetition rate goes up. A related measurement is reported in [150] for which pulsing was paused for intervals of tens of seconds to the order of a day. An enhanced breakdown rate was observed following the pause, with a greater enhancement for longer pauses. The explanation given in [150] for this effect is that gas is adsorption during the field off interval between pulses and during pauses. The longer the time between pulses, either through a lower repetition rate or through inhibited pulsing, more residual gas such as water is re-adsorbed. This adsorbed gas then results in an increased probability of breakdown. This highlights that there are multiple breakdown trigger mechanism.
@@
Finally, a dependence of breakdown rate that has been observed in DC systems is on the distance between electrodes. Electric field is usually calculated as where is the applied voltage. It has been observed that most experimental results obtained at various electrode separations scale less than linearly with the gap, as . This is illustrated in Fig. 26. This phenomenon has already been identified in the literature both in dc [106, 4] and at high frequency [102] as being an implicit effect of the applied voltage. A deeper understanding of its origin will be discussed in more detail in Section IX.
@@ III.6 Dependence on material
There are preferred metals for different high-field applications, for example, copper in radio frequency devices, stainless steel in large high-voltage systems as well as focusing elements in scanning electron microscopes, and alloys of copper in circuit breakers and X-ray tubes. These have usually been chosen through long experience and trial-and-error testing. The demand for a very high accelerating gradient in linear colliders has motivated a renewed look at the field holding capabilities of different materials to see if a better-performing metal than copper could be found. In addition, comparisons of the results of systematic tests of different materials have yielded fundamental insights into the breakdown mechanism. The material tests carried out by the linear collider and related collaborations and summarized here include those carried out using RF accelerating structures and using both the tip-to-sample and pulsed DC systems.
@@
Tests of the field holding capability of copper, molybdenum and tungsten have been carried out using 30 GHz and 11 GHz accelerating structures using both clamped irises and quadrant type structures. The results show that in certain tests, molybdenum and tungsten reached higher accelerating gradients than copper. However, the improvement was not seen in all cases, which is likely due to the difficulties of machining the refractory materials to the shapes required for RF structures, in particular resulting in a poor surface finish, as well as the difficulty in incorporating the materials in RF structures. In addition, the sometimes observed field holding improvement were not sufficient overall to overcome the poorer electrical and thermal conductivities. These results are described in [186, 183, 39].
@@
Compared to radio frequency structures, dc systems are better adapted for comparative material tests since electrodes have simpler geometries and thus are easier to fabricate, and electrodes do not have current carrying joints. In addition, low electrical conductivity materials can be tested in the dc system.
@@
The anode tip system has been used to test the maximum field holding capacity of a large number (16) of materials. The materials show a large spread in maximum field holding strength, as shown in Fig. 28. The test have yielded the insight that pure metals are ranked by their crystal structure with FCC, BCC and HCP holding progressively higher surface fields. The crystal structures also have progressively higher energy barriers to dislocation movement, providing evidence that dislocations play a major role in breakdown initiation. Cobalt, a relatively soft and low melting point material but with an HCP crystal structure, was only tested once the crystal structure dependence was observed. It held a high field as predicted. The results are reported in [35].
@@
Samples of different type in terms of material, surface finishing, heat treatment, etc. have been tested with the anode tip system, [76, 34]. The test protocol of these test was to apply a progressively increasing voltage until a breakdown occurred, giving the threshold field. Voltage was then reduced to the staring value, and the cycle restarted until next breakdown. A steady state, with some level of fluctuation, is reached after few or several breakdowns, depending on the material. This progressive improvement process is defined as conditioning, as will be discussed further in III.4. The breakdown field varies from pulse to pulse but an average maximum breakdown field can be computed. This type of experiment has been performed for many sample-tip pairs of the same material and reported in several publications [76, 34, 35]. One particular experiment has involved heterogeneous cathode-anode pairs made of two different materials, namely Ti and W. In this case, it was shown that the maximum breakdown field was equal the one of the cathode, when tested in a same-material cathode-anode pair [34]. This supports observations that breakdown starts on the cathode and that the maximum field is determined by the cathode material, for geometrically identical configurations.
@@
The pulsed dc system has been used to test the field holding properties of a range of materials. Compared to the anode tip tests, the pulsed dc test focused more on alloys of copper rather than pure metals. This choice was motivated by the high electrical conductivity requirement for potential use in a radio frequency structure. In addition, some materials with a high solubility of hydrogen were tested for use in radio frequency quadrupoles, a linear accelerator device used to accelerate and bunch low energy particles such as protons and ions, which are sometimes subject to H- beam loss [155]. The results are shown in Fig. 29. In this case the electrodes were tested in low breakdown rate conditions, in contrast to the maximum field holding mode of tests done in the anode tip system, which were effectively carried out at a breakdown rate equal to 1. The smaller range of surface field may be due to the choice of materials and possibly to the low breakdown rate tests as opposed to maximum field holding tests, a divergence also suggested by Fig. 24.
@@
There have also been systematic tests of single-cell RF cavities made from different materials. These have the same design as those used for the cryogenic tests described above and reported in [21]. These have included cavities built from different copper alloys and different hardness states, and some results appear in Fig. 24. The general conclusion of the group is that harder materials can hold higher gradients. In the case of the comparison of heat-treated and not heat-treated copper, this result can be compared to the conditioning results shown in Fig. 19 where hard copper conditions much more quickly than soft copper, but both arrive at the same limiting value. A further publication from this group is [41].
@@ III.7 Field emission
Electron field emission is another phenomenon that occurs when high surface electric fields are applied to conducting surfaces. In this section we will present experimental measurements of field emission in both RF and DC conditions. There are both theoretical and practical motivations for measuring and understanding field emission. An example of a practical consideration is found in the field of high-gradient accelerators where field emission can interfere with operation. The problem occurs when field emitted currents, often referred to as dark currents, are captured by RF fields and accelerated over longer distances inside the accelerator. These accelerated currents can create an unwanted background signal for beam instrumentation and can also lead to X-ray radiation and material activation when accelerated dark currents hit accelerator components.
@@
The measurements of field emitted currents presented here were made to contribute to understanding the breakdown initiation process. One aspect has been to provide experimental data to complement the simulation studies of the direct role of field emission in the breakdown initiation as described in Sec. VI. These simulations study breakdown initiation through localized heating of field emission enhancing features on cathodic surfaces. Another aspect has been to try to provide direct experimental evidence of the fluctuations that are predicted in the analysis of dislocation dynamics described in Sec. IV. In addition an attempt has been made to study field emission current signals using machine learning techniques to see if a breakdown precursor could be identified. Although inconclusive, such a study is described in [131].
@@
The radio frequency field emission measurements presented here were taken in pulsed mode, just as the breakdown measurements described throughout this review. The dc field emission measurements were taken with continuously applied fields, whereas the breakdown measurements were taken in pulsed mode. These two modes of the pulsed dc system are described in in Subsec. II.3. One typical type of measurement is the , current vs. voltage curve. This is a widely spread type of measurement.
@@
The wide variety of conditions in which breakdowns have been studied have given opportunities for investigation of field emission as well, and new perspectives on the relationship of the two processes. An example of such a measurement is a comparison of the field emission behavior in the same system for different materials, as shown in the upper plot of Fig. 30. A striking feature of this plot is that, at equivalent field, materials that reached higher fields showed much lower current compared to those that achieved lower fields. The field emission measurements of the different materials we made in conjunction with the field holding capability measurements of the same materials in order to better understand how the two processes are related.
@@
data can also be plotted in Fowler-Nordheim and Murphy-Goode forms in order to extract the so-called field enhancement factor . A field enhancement factor is a way of matching measured field emission current to the applied field dependence predicted by the Fowler-Nordheim equation, 3. This correction factor generally falls in the range from 50 to a few hundred. To the knowledge of the authors, such large correction factors are required for all field emission measurements made on macroscopic systems. The only exception where a correction factor is not required is for nanometric-scale tips, for example [187]. The correction factor is of the order of 30 to 60 for the high-gradient accelerating structures described in this review, and of the order of 100 to 200 for the electrodes used in the pulsed DC system. The data shown in the upper part of Fig. 30 is re-plotted in Fowler-Nordheim form, along with extracted field enhancement values is shown in the lower part of Fig. 30. The field enhancement factor is discussed further from a theoretical point of view in Sec. V.2.
@@
It should be noted that in some cases high values recorded typically at the beginning of an experiment can usually be attributed to surface contamination [132]. These may be removed with some surface treatments [89], or are more effectively destroyed upon first breakdowns, resulting in surface conditioning. However, the decrease in the field enhancement factor observed when a device is conditioned is not only linked to extrinsic properties like contamination, but has some more deep links with intrinsic material properties. An example of this was seen during the high-power test of a 3 GHz backward traveling wave accelerating structure as described in [108]. Here the field enhancement factor decreased from 76 to 46 over the conditioning process as shown in Fig. 31. In this test remained relatively constant over the conditioning process. Another example when remained constant during operation was in a test with the anode tip system as described in [33] and shown in Fig. 32. In this experiment the system was operated in a breakdown threshold mode, where on each pulse the voltage was ramped up until breakdown occurred. The plot shows that the field enhancement factor measured between pulses was an accurate predictor of the field at which the next breakdown would occur through the relationship . This confirms earlier measurements [4, 20, 83]. Other examples showing the evolution of during conditioning can be found in [12, 178]. These particular experiments indicate that there is an intrinsic material property that determines the local breakdown field threshold . Other experiments have not shown the same relationship.
@@
Another experimental condition where higher field holding is associated with lower field emission is for cryogenically cooled electrodes as described in [61]. Fig. 25 shows a significant increase in field for a given breakdown rate when electrodes are cooled to cryogenic temperatures, as discussed in Subsec. III.5. Field emission data from the same electrodes show, for an equivalent applied field, a significant drop in field emission current when cooled. On the other hand, the field emission current becomes much higher than at room temperature at the heightened field levels made possible by the cyrogenic cooling. In addition, the field emission becomes much more stable than at room temperature, resulting in close to ideal straight lines in the Fowler-Nordheim plots shown in Fig. 25 over a wide range of field emission current. The two latter observations indicate that field emission sites become extremely stable at cryogenic temperatures.
@@
Another important aspect of field emission is that light is emitted. Measurements of the light associated with field emission are described in [82, 137]. The intensity of emitted light is closely correlated with emitted current as shown in 34. The origin of the light is not yet certain, but it is currently believed to be from optical transition radiation emitted when field emitted currents strike the anode. Such a mechanism would give the direct proportionality of light and current that is observed, compared to, for example, black body radiation from the emitter site which would give a more complex relationship. Optical emission opens new ways of monitoring field emission since it avoids potential damage to electronics which can occur due to the multi-order-of-magnitude current surge that occurs if a breakdown occurs during a field emission measurement. Mitigating electronic measures, such as series resistances can interfere with the measurement itself. In addition, the emitted light has a very large bandwidth especially compared to electronic current measurements. This may improve the fluctuation measurements described above.

### v2

@@ III.5 Dependencies of breakdown rate
Breakdown rate depends very strongly on field level, pulse length and temperature. These dependencies are very important for device operation, and are essential input for the dislocation model described in Sec. IV.
The breakdown rate of a high-field system exhibits a very strong dependence on applied field level. They are an important input for the dislocation models described in Sec. IV.In both pulsed rf and dc systems, this dependence is typically,
This has been reported for both radio frequency [61] and pulsed dc systems [166]. Example of breakdown rate vs. gradient measurements are shown in Figs. 22 and 23. Breakdown rate data for radio frequency experiments is typically presented in units of , with the values computed by dividing the device breakdown rate by the active length of the structure. This is done so that the breakdown rate of a single accelerating structure can be extrapolated to a linac containing many structures. Breakdown rate data from the pulsed dc system is typically presented for the device, without normalizing for example for the surface area of the electrode.
@@
The breakdown rate also displays a dependence on pulse length. In the traveling wave rf accelerating structures tested in the linear collider programs, measured breakdown rates typically have shown a dependency of,
@@
where is the pulse length, as for example reported in [61]. An example where this dependency can be observed is by comparing the shapes of the conditioning curves shown in Figs. 18 and 19. In both cases conditioning has been carried out at successively longer pulse lengths. The change of pulse length can be clearly seen in Fig. 18. In Fig. 19 the electric field is normalized by , the conditioning curve becomes smooth showing that the pulse length scaling is correct. It is important to note that these pulse lengths for the accelerating structure tests ranged from about 50 to 300 ns, with the minimum time give by the filling time of the structures (tests were rarely carried out below the filling time) and the maximum time is given by the power limit of the klystron-based power sources driving the structures. Breakdown rate as a function of pulse length measurements have also been carried out using the pulsed dc system, but here pulse lengths were in the range of around 500 ns up to 1 ms. In this system and pulse length range the pulse length dependence was indiscernible as shown in Fig. 24. It appears that the pulse length dependence saturates at around 500 ns. The mechanism of pulse length dependence is discussed further in subsection IV.5. Finally, breakdowns can even occur for pulses that short compared to the plasma formation time, 10 ns and less. Very short pulse breakdowns are not covered further in this report but data is available in [20, 32, 170].
@@
An important insight into the character of the breakdown mechanism has been found when combining the observation that the breakdown timing distribution is flat, as shown in Fig. 17 and discussed in subsection III.3 and even while a strong pulse length dependence is observed in rf tests. Since the breakdown rate dependence is not produced by a progressively greater density of breakdowns at the end of the pulse, the conclusion must be that the system displays a memory of the pulse length of previous pulses. This is analogous to a fatigue process where a material will tend to fail after a certain number of stress and de-stress cycles, but fails at a random moment within a cycle. This memory effect is an important motivation for the dislocation-based breakdown initiation model described in Sec. IV.
@@
Another observed dependence of breakdown rate is on temperature. For this dependence, the breakdown rate at a given field level goes down significantly when a high-field system is cooled to cryogenic temperatures. Alternatively one can say that system supports a higher field for a given breakdown rate when cooled. This type of measurement has required construction of specialized versions of the rf and pulsed dc systems.
@@
An early high-field measurement of the effect of cooling on field was carried out in the CTF2 facility using a beam-driven cavity cooled to liquid nitrogen temperature [20]. This experiment did not show any temperature dependence, however only the maximum field holding level was measured, that is the field level where the breakdown rate approaches 1. A subsequent rf experiment was carried out using a cryogenically cooled cavity as described in [22], in which the effect of temperature on field and breakdown rate was measured at low, range, breakdown rates. The experiment used an rf test stand similar to the type described in Sec. II but where the cavity was operated at cryogenic temperatures using a cryo-cooler. When cooled to 45 K, the cavity showed a very clear, approximately 20, increase of field level for a breakdown rate of approximately . A plot of the effect of temperature on gradient is shown in Fig. 25. The publication also reports that field emission at the increased field levels enabled by the cooling was strong enough to result in a measurable dynamic decrease of cavity .
@@
A similar increase of field for a given breakdown rate at lowered temperatures has also been observed in pulsed dc experiments [64]. In this case, the experimental setup consisted of a pulsed dc system similar to that described in Sec.II but where the electrodes were cooled to cryogenic temperatures. Here also the field increased by approximately 20 for a normalized breakdown rate when cooled to a temperature of around 60 K and conditioned, with further improvement as the system was cooled further to around 30 K. Data showing the increase of field for a normalized breakdown rate in the pulsed dc system is shown in Fig. 26. The dc systems also showed enhanced field emission at the elevated field levels, and we will return to this in subsection III.7.
@@
Another dependence of breakdown rate that has been observed is on repetition rate and more generally, the time interval between pulses. In a pulsed dc experiment described in [159], the breakdown rate was measured as the repetition rate was varied from 10 Hz to 6 kHz and is shown in Fig. 27. The breakdown rate decreases as the repetition rate increases. A related measurement is reported in [159] for which pulsing was paused for intervals between tens of seconds to the order of a day. An enhanced breakdown rate was observed following the pause, with a greater enhancement for longer pauses. The explanation given in [159] for this effect is that gas is adsorbed during the field-off interval between pulses and during pauses. The longer the time between pulses, either through a lower repetition rate or through inhibited pulsing, the more residual gas, such as water vapor, is re-adsorbed. This adsorbed gas then results in an increased probability of breakdown.
@@
Finally, we discuss the dependence of the breakdown rate that has been observed in dc systems on the distance between electrodes. The surface electric field is , where is the applied voltage. It has been observed that most experimental results obtained at various electrode separations the maximum voltage scales less than linearly with the gap, as . Thus the maximum surface electric field decreases for increasing inter-electrode gap. This is illustrated in Fig. 28. This phenomenon has already been identified in the literature, both in dc [114, 4] and at high frequency [109] as being an implicit effect of the applied voltage. The theory which also explains the gap dependence will be developed in Section IX.
@@ III.6 Dependence on material
There are preferred metals for different high-field applications, for example, copper in radio frequency devices, stainless steel in large high-voltage systems as well as focusing elements in scanning electron microscopes, and alloys of copper in circuit breakers and X-ray tubes. These have usually been chosen through long experience and trial-and-error testing. The demand for a very high accelerating gradient in linear colliders motivated a renewed look at the field holding capabilities of different materials to see if a better performing metal than copper could be found. In addition, comparisons of the results of systematic tests of different materials have yielded fundamental insights into the breakdown mechanism. The material tests carried out by the linear collider and related collaborations summarized here include those carried out using rf accelerating structures and using both the anode tip and pulsed dc systems.
@@
Tests of the field holding capability of copper, molybdenum and tungsten have been carried out using Ka and X-band accelerating structures using both clamped irises and quadrant type structures. The results show that in certain tests, molybdenum and tungsten reached higher accelerating gradients than copper. However, the improvement was not seen in all cases, which is likely due to the difficulties of machining the refractory materials to the shapes and surface finish required for rf structures as well as the difficulty in joining parts. In addition, the sometimes observed field holding improvement were not sufficient overall to overcome the poorer electrical and thermal conductivities. These results are described in [196, 193, 42].
@@
Compared to radio frequency structures, dc systems are better adapted for comparative material tests since electrodes have simpler geometries, and thus are easier to fabricate and do not have current carrying joints. In addition, low electrical conductivity materials can be tested in the dc system.
@@
The anode tip system has been used to test the maximum field holding capacity of a large number (16) of materials, resulting from averaging several ramp-to-breakdown experiments. The materials show a large spread in maximum field holding strength, as shown in Fig. 29. The tests yielded the insight that pure metals are ranked by their crystal structure with FCC, BCC and HCP holding progressively higher surface fields. These crystal structures also have progressively higher energy barriers to dislocation movement, providing evidence that dislocations play a major role in breakdown initiation. Cobalt, a relatively soft and low melting point material but with an HCP crystal structure, was only tested once the crystal structure dependence was observed. It held a high field as predicted. The results are reported in [38]. This was an important step in the development of the dislocation model as described in Sec. IV Plastic Material Response.
@@
Samples of different type in terms of material, surface finishing, heat treatment, etc. have been tested with the anode tip system, [80, 37]. The ramp-to-breakdown test protocol was to apply a progressively increasing voltage until a breakdown occurred, giving the threshold field. Voltage was then reduced to the initial value, and the cycle restarted until next breakdown. A steady state, with some level of fluctuation, was reached after few or several breakdowns, depending on the material. The breakdown field varies from pulse to pulse but an average maximum field holding capacity could be computed. This type of experiment has been performed for many sample-tip pairs of the same material and reported in several publications [80, 37, 38]. One particular experiment involved cathode-anode pairs made of two different materials, namely Ti and W. In this case, the maximum breakdown field corresponded to that of the cathode material when tested in a same-material cathode-anode pair [37]. This supports observations that breakdown starts on the cathode and that the maximum field is determined by the cathode material.
@@
The pulsed dc system has been used to test the field holding properties of a range of materials. Compared to the anode tip tests, the pulsed dc test focused more on alloys of copper rather than pure metals. This choice was motivated by the high electrical conductivity requirement for potential use in a radio frequency structure. In addition, some materials with a high solubility of hydrogen were tested for use in radio frequency quadrupoles, a linear accelerator device used to accelerate and bunch low energy particles such as protons and ions, which are sometimes subject to H- beam loss [164]. The results are shown in Fig. 30. These results are for low breakdown rate operation performed with large electrodes, while those shown in Fig. 29 are for ramp-to-breakdown operation performed with the anode tip system, as described in Sec. II.3. In ramp-to-breakdown operation the voltage is raised until a breakdown occurs, giving an effective breakdown rate of 1. The smaller range of surface field shown in Fig. 30 may be due to the choice of materials and possibly to the difference between low breakdown rate and ramp-to-breakdown testing.
@@
There have also been systematic tests of single-cell rf cavities made from different materials. These have the same design as those used for the cryogenic tests described above and reported in [22]. These have included cavities built from different copper alloys and different hardness states, and some results appear in Fig. 25. The general conclusion is that harder materials hold higher gradients. In the case of the comparison of heat-treated and not heat-treated copper, this result can be compared to the conditioning results shown in Fig. 20 where hard copper conditions much more quickly than soft copper, but both arrive at the same limiting value. A further publication is [44].
@@ III.7 Field emission
Electron field emission is another phenomenon that occurs when high surface electric fields are applied to conducting surfaces. In this section, we will present experimental measurements of field emission in both rf and dc conditions. There are both theoretical and practical motivations for measuring and understanding field emission. An example of a practical consideration is found in the field of high-gradient accelerators, where field emission can interfere with operation. The problem occurs when field-emitted electrons, often referred to as dark currents, are captured by rf fields and accelerated over long distances inside the accelerator. These accelerated electrons can create an unwanted background signal for beam instrumentation and can also lead to X-ray radiation and material activation when they hit accelerator components.
@@
The measurements of field-emitted currents presented here were made to contribute to understanding the breakdown initiation process. One aspect has been to provide experimental data to complement the simulation studies of the direct role of field emission in the breakdown initiation as described in Sec. VI. These simulations study breakdown initiation through localized heating of field emission enhancing features on cathodic surfaces. Another aspect has been to try to provide direct experimental evidence of the fluctuations that are predicted in the analysis of dislocation dynamics described in Sec. IV. In addition, an attempt has been made to study field emission current signals using machine learning techniques to see if a breakdown precursor could be identified. Although inconclusive, such a study is described in [140].
@@
The radio frequency field emission measurements presented here were taken in pulsed mode, just as the breakdown measurements described throughout this review. The dc field emission measurements were taken with continuously applied fields, whereas the breakdown measurements were taken in pulsed mode. These two modes of the pulsed dc system are described in subseciton II.3. One typical type of measurement is the , current vs. voltage curve.
@@
The wide variety of conditions in which breakdowns have been studied have given opportunities for investigation of field emission as well, and new perspectives on the relationship of the two processes. An example of such a measurement is a comparison of the field emission behavior in the same system for different materials, as shown in the upper plot of Fig. 31. A striking feature of this plot is that, at equivalent field, materials that reached higher fields showed much lower current compared to those that achieved lower fields. Field emission and field holding measurements of different materials were made to better understand how the two processes are related.
@@
data can also be plotted in Fowler-Nordheim and Murphy-Goode forms in order to extract the so-called field enhancement factor . A field enhancement factor is a way of matching measured field emission current to the applied field dependence predicted by the Fowler-Nordheim [55]. This correction factor generally falls in the range from 50 to a few hundred. To the knowledge of the authors, such large correction factors are required for all field emission measurements made on macroscopic systems. The only exception where a correction factor is not required is for nanometric-scale tips, for example [197]. The correction factor is of the order of 30 to 60 for the high-gradient accelerating structures described in this review, and of the order of 100 to 200 for the electrodes used in the pulsed dc system. The data shown in the upper part of Fig. 31 is re-plotted in Fowler-Nordheim form, along with extracted field enhancement values in the lower part of Fig. 31. The field enhancement factor is discussed further from a theoretical point of view in Sec. VI.
@@
It should be noted that in some cases, high values recorded at the beginning of an experiment can usually be attributed to surface contamination [141]. These may be removed with some surface treatments [94], or are more effectively destroyed upon first breakdowns, resulting in surface conditioning. However, the decrease in the field enhancement factor observed when a device is conditioned is not only linked to extrinsic properties like contamination, but has some more deep links with intrinsic material properties. An example of this was seen during the high-power test of a 3 GHz backward traveling wave accelerating structure as described in [116]. Here the field enhancement factor decreased from 76 to 46 over the conditioning process as shown in Fig. 32. In this test remained relatively constant over the conditioning process. Another example where remained constant during operation was in a test with the anode tip system as described in [36] and shown in Fig. 33. In this experiment the system was operated in a breakdown threshold mode, where on each pulse the voltage was kept at the same level, until breakdown occurred. The plot shows that the field enhancement factor measured between pulses was an accurate predictor of the field at which the next breakdown would occur through the relationship . This confirms earlier measurements [4, 21, 88]. Other examples showing the evolution of during conditioning can be found in [12, 188]. These particular experiments indicate that there is an intrinsic material property that determines the local breakdown field threshold . Other experiments have not shown the same relationship.
@@
Lower field emission for an equivalent field level is also observed with cryogenically cooled electrodes as described in [64]. We have seen in Fig. 26 of Sec. III.5 a significant increase in field for a given breakdown rate when electrodes are cooled to cryogenic temperatures. Field emission data from the same electrodes show that, for an equivalent applied field, a significant drop in field emission current when cooled. On the other hand, the field emission current becomes much higher than at room temperature at the heightened field levels only accessible when the electrodes are at cryogenic temperatures. This is because the field emission current level rises exponentially, which dominates over the reduced field emission at lower fields. This effect was also observed in the rf measurements described in [22] in which a dynamic loss of cavity quality factor due to field emitted current beam loading was observed. In addition, the field emission becomes much more stable at cyrogenic temperatures than at room temperature, resulting in close to ideal straight lines in the Fowler-Nordheim plots shown in Fig. 34 over a wide range of field emission current. These cryogenic measurements indicates that field emission sites become extremely stable at low temperatures.
@@
Another important aspect of field emission is light is emission. Measurements of the light associated with field emission are described in [87, 146]. The intensity of emitted light is closely correlated with emitted current, as shown in Fig. 35. The origin of the light is not yet certain, but it is currently believed to be from optical transition radiation emitted when field emitted currents strike the anode. Such a mechanism would explain the observed direct proportionality of light and current compared to, for example, black body radiation from the emitter site which would give a more complex relationship. Optical emission opens new ways of monitoring field emission since it avoids potential damage to electronics which can occur due to the multi-order-of-magnitude current surge that occurs if a breakdown occurs during a field emission measurement. Mitigating electronic measures, such as using series resistances can interfere with the measurement itself. In addition, the emitted light has a very large bandwidth especially compared to electronic current measurements. This may improve the fluctuation measurements described above.